I have pure silver 1300 gms. I have mixture(x) 65% silver( + 35 % other metals).How many gms of (x) to be added with pure silver (1300 gms) to get 73 % of new mixture?
Answer by richwmiller(17219)

We have 1300 gms of 100% silver
We want to add some 65% silver to make 73% silver
0.65*x+1*1300=0.73(1300+x)
0.65*x+1300=949+0.73x
0.65x-0.73x=949-1300
-0.08x=-351
x=4387.5 gms at 65% (2851.875) is added to 1300 gms at 100%(1300) to make 73% of 5687.5 gms (4151.875)
check
0.65*4387.5+1*1300=0.73(1300+4387.5)
2851.875+1300=0.73(5687.5)
4151.875=4151.875
ok
codemixadd